OpenStudy (erich):

Solve for X (a-bx)/c = 1/2c + px/c

OpenStudy (erich):

a-bx-px / c = 1 / 2c

OpenStudy (anonymous):

a-bx= 1/2 + px (multiply the entire exuation by c) -bx -px = 1/2 -a (isolate the x terms) bx + px = a - 1/2 (multiply by -1 for simplicity) (b+p)x = a- 1/2 x= (a - 1/2)/(b+p)